OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 13 and a13 = –59?

OpenStudy (anonymous):

hi guyz

OpenStudy (apoorvk):

okay so we know nth term of an AP is a1 + (n-1)d .. here d is the common difference between each term. so, a1=13.... (i) a1 + (13-1)d=-59 ...(ii) Put value of a1 from (i) in (ii) to solve it and get 'd' then find 32nd term by the formula for any nth term!! post what you get and verify, or if you face any problems!

OpenStudy (anonymous):

@kmila03 Can you follow apoorvk's instruction?

OpenStudy (anonymous):

mmhm not really im trying ...

OpenStudy (anonymous):

a_13 = a_1 + 12d Do you know this formula by now?

OpenStudy (anonymous):

actually thats the problem i never seen this before

OpenStudy (anonymous):

Memorize this formula, right now just plug number in!

OpenStudy (anonymous):

ok, but what is 12d?

OpenStudy (anonymous):

d is what we're looking for d = ...

OpenStudy (apoorvk):

d is the variable.. the common difference any two successive terms of an AP that you gotta find!

OpenStudy (anonymous):

12d = a_13 - a_1 =>d = (a_13 - a_1) / 12 Can you plug in the numbers now?

OpenStudy (anonymous):

so, can you guys do it. so ill see the steps?

OpenStudy (apoorvk):

okay i am continuing from my first post above..... just tell me do you understand linear equations?? did you understand the equations i wrote? and do you understand the question?

OpenStudy (anonymous):

yes... so if i plug in the number from =>d = (a_13 - a_1) / 12 is = d=-6 ?

OpenStudy (apoorvk):

that's it kmila!!!!! you got it!!! what was there to be afraid of?? :) now can you find the 32nd term??

OpenStudy (anonymous):

thanks :)..... but how can i find 32nd?

OpenStudy (apoorvk):

okay any nth term "Tn" of an AP is Tn=a + (n-1)d where 'a' is the first term and 'd' is the common diff. so, n is 32 here, first term is given and you just found out d. plug in the values and get the answer!

OpenStudy (apoorvk):

it may so happen in future questions that the first term may not be given straightaway, but any two terms say 4th and 6th are provided, in that case you need to form 2 linear equations in 'a' and 'd' and solve to find 'a' and 'd'. then proceed with what the question asks!

OpenStudy (anonymous):

so it will be -173?

OpenStudy (anonymous):

Yeehhh, you got it :D

OpenStudy (apoorvk):

Riiighttttt-O!!!!!!! there you are!!

OpenStudy (anonymous):

ok, thank u guys so much!!!

OpenStudy (anonymous):

We hope you learn the formula AND keep PRACTICE :)